# LoRA Done RITE: Robust Invariant Transformation Equilibration for LoRA Optimization

### v1

@@ 3.3 Theoretical Analysis
Following previous work (Gupta et al., 2018; Feinberg et al., 2023), we provide a convergence analysis of the proposed algorithm within the online optimization framework (Hazan et al., 2016; Shalev-Shwartz et al., 2012).
In online convex optimization setting, a parameter ${\bm{\theta}}_{t}\in\mathcal{K}$ is chosen iteratively, where $\mathcal{K}$ is a convex decision set. After each decision- ${\bm{\theta}}_{t}$, a convex loss function $f_{t}$ is revealed, potentially chosen adversarially. The regret accumulated by the algorithm up to step $T$ is defined as
@@
We assume for the fine-tuned weight ${\bm{Z}}$ of each layer, the convex decision set imposes the following constrains:
$\|{\bm{A}}\|_{F}\leq D_{\bm{A}},\|{\bm{B}}\|_{F}\leq D_{\bm{B}},$
@@
Let
$\bar{{\bm{X}}}_{{\bm{A}}_{t}}=(\bar{{\bm{V}}}_{{\bm{A}}_{t}}+\rho_{{\bm{A}}_{t}}{\bm{I}})^{-1/2}$
be the unmagnified preconditioner ${\bm{P}}_{{\bm{A}}_{t}}={\bm{U}}_{{\bm{B}}_{t}}^{\top}{\bm{U}}_{{\bm{B}}_{t-1}}$, and ${\bm{Q}}_{{\bm{A}}_{t}}={\bm{R}}_{{\bm{B}}_{t}}^{-\top}{\bm{R}}_{{\bm{B}}_{t-1}}^{\top}$, then we have
$\|\bar{{\bm{X}}}_{{\bm{A}}_{t}}^{-1}-{\bm{Q}}_{{\bm{A}}_{t}}\bar{{\bm{X}}}_{{\bm{A}}_{t-1}}^{-1}{\bm{Q}}_{{\bm{A}}_{t}}^{\top}\|\leq\mu\|\bar{{\bm{X}}}_{{\bm{A}}_{t}}^{-1}-{\bm{P}}_{{\bm{A}}_{t}}\bar{{\bm{X}}}_{{\bm{A}}_{t-1}}^{-1}{\bm{P}}_{{\bm{A}}_{t}}^{\top}\|.$
This assumption essentially constrains the change in ${\bm{R}}_{{\bm{B}}_{t}}$ to be relatively smooth. Under this assumption, we can establish the following stronger convergence result:
@@
$\frac{1}{T}\sum\nolimits_{t=1}^{T}\nabla{\bm{\theta}}_{t}^{\top}({\bm{\theta}}_{t}-{\bm{\theta}}^{*})=\mathcal{O}(GD_{\bm{A}}D_{\bm{B}}T^{-1/2}).$
@@ 4 Related Work
LARS (You et al., 2017) and Lamb (You et al., 2020a) are layer-wise adaptive optimization methods originally designed for large batch training. They dynamically adjust the update norm for each weight matrix based on its current norm, which ensures scalar scale invariance. Nonetheless, they still lack transformation invariance.
@@
Gemma-2B Adam 83.76 45.31 24.26 64.00 54.33
LoRA+ 83.75 45.31 23.65 64.40 54.28
ScaledAdam 83.52 45.22 23.96 64.80 54.38
Shampoo 83.26 44.88 23.35 63.60 53.77
Lamb 86.60 47.35 26.76 68.00 57.18
LoRA-RITE 87.28 49.06 30.10 68.80 58.81
@@ 5 Experimental Results
We evaluated the proposed LoRA optimizer against other optimizers across a range of datasets. This included the Super-Natural Instructions dataset, a comprehensive collection of diverse NLP tasks, as well as four standard LLM benchmarking datasets.
@@
  • Lamb (You et al., 2020b): A variant of Adam that normalizes the updates for each layer based on the norm of the parameters.
@@
For each optimizer applied on each data set, we search for the best learning rate from $2*10^{-6}$ to $2*10^{-2}$. The other hyperparameters are listed in the appendix. For most of the experiments we chose rank $r=16$ for LoRA, based on the ablation study over the rank at the end of experiments. We conduct experiments on Gemma (Gemma Team et al., 2024)  2B, 7B, and mT5-XXL (Xue et al., 2021) using TPUs.
@@
We conduct an ablation study on the choice of different LoRA ranks and model architectures. Specifically, we considered rank 4 and 16 on both Gemma 2B (decoder only) and mT5-XXL (encoder-decoder) on the OpenBookQA dataset. As we can see from Table 3, our proposed method performs consistently well across different LoRA ranks. Furthermore, our method can be successfully applied to mT5-XXL which has an encoder-decoder architecture, showing the generalizability of the proposed optimizer. We also show the experiment results of Gemma2 in the appendix.
@@
Table 4: Number of training steps per second for different optimizers. LoRA-RITE has a small overhead compared with first-order methods.
@@ 6 Conclusions
Although this work introduces a better optimizer for LoRA, it is important to acknowledge that LoRA itself has limitations. For instance, LoRA has smaller representational power and may result in a small performance decrease compared to full fine-tuning. Also, how to select rank to strike a good trade-off between efficiency and accuracy may be non-trivial in practice.
@@ A.3 Proof of Theorem 1
Let $\|{\bm{A}}_{1}\|={\bm{\theta}}(n^{a})$, $\|{\bm{B}}_{1}\|={\bm{\theta}}(n^{b})$, $\|\nabla{\bm{Z}}\|={\bm{\theta}}(n^{c})$, $\eta={\bm{\theta}}(n^{d})$, where $\eta$ is the learning rate and $n$ is the network width. Since ${\bm{Z}}={\bm{A}}_{1}{\bm{B}}_{1}^{\top}$, from chain rule we know $\nabla{\bm{A}}=\nabla{\bm{Z}}{\bm{B}}$ and $\nabla{\bm{B}}=\nabla{\bm{Z}}^{\top}{\bm{A}}$. Since the update rule is symmetric, we can express the updates as
@@
where ${\bm{x}}$ is the input vector, by selecting a proper learning rate $\eta={\bm{\theta}}(n^{d})$.
@@ A.4 Proof of Theorem 2
For matrix ${\bm{X}}_{{\bm{A}}}\in\mathbb{R}^{m\times r}$, ${\bm{H}}_{{\bm{A}}}\in\mathbb{R}^{r\times r}$, we call them consistent if
${\bm{X}}_{{\bm{A}}}{\bm{U}}_{{\bm{B}}}^{\top}\in\mathbb{R}^{m\times n}$
and
${\bm{U}}_{{\bm{B}}}{\bm{H}}_{{\bm{A}}}{\bm{U}}_{{\bm{B}}}^{\top}\in\mathbb{R}^{n\times n}$
are respectively the same across all equivalent LoRA pairs.
@@
First, one should note the fact that
${\bm{U}}_{\bm{B}}{\bm{U}}_{\bm{B}}^{\top}$
is the same across all equivalent pairs. Thus,
${\bm{U}}_{\bm{B}}(\bar{\nabla}{{\bm{A}}})^{\top}\bar{\nabla}{{\bm{A}}}{\bm{U}}_{\bm{B}}^{\top}={\bm{U}}_{\bm{B}}{\bm{U}}_{\bm{B}}^{\top}\nabla{\bm{Z}}^{\top}\nabla{\bm{Z}}{\bm{U}}_{\bm{B}}{\bm{U}}_{\bm{B}}^{\top}$
implies (¯$(\bar{\nabla}{{\bm{A}}})^{\top}\bar{\nabla}{{\bm{A}}}$ is consistent.
This combined with the fact that ${{\bm{P}}_{{\bm{A}}_{t}}}\bar{{\bm{V}}}_{{\bm{A}}_{t-1}}{\bm{P}}_{{\bm{A}}_{t}}^{\top}$ is consistent if $\bar{{\bm{V}}}_{{\bm{A}}_{t-1}}$ is consistent and that $\bar{{\bm{V}}}_{{\bm{A}}_{0}}=\bf{0}$ implies $\bar{{\bm{V}}}_{{\bm{A}}_{t}}$ is consistent.
@@
both $\bar{{\bm{S}}}_{{\bm{A}}_{t}}$ and $\bar{{\bm{M}}}_{{\bm{A}}_{t}}$ are consistent, which completes our proof.
@@ A.5 Proof of Theorem 3
$\bar{{\bm{X}}}_{{\bm{A}}_{t}}=\bar{{\bm{V}}}_{{\bm{A}}_{t}}^{-1/2}.$
Then by Lemma 2, for the ${\bm{A}}$ factor, we have
$\sum_{t=1}^{T}\operatorname{vec}(\nabla{\bm{A}}_{t})^{\top}\operatorname{vec}(\delta{\bm{A}}_{t})=\sum_{t=1}^{T}\operatorname*{Tr}(\nabla{\bm{A}}_{t}^{\top}\delta{\bm{A}}_{t})$ (19)
$=$ $\eta\sum_{t=1}^{T}\|\nabla{\bm{A}}_{t}\|_{{\bm{X}}_{{\bm{A}}_{t}}}^{2}=\eta\sum_{t=1}^{T}\|\bar{\nabla}{\bm{A}}_{t}\|_{\bar{{\bm{X}}}_{{\bm{A}}_{t}}}^{2}\leq 2\eta\operatorname*{Tr}(\bar{{\bm{V}}}_{{\bm{A}}_{T}}^{1/2}).$
Since
$\operatorname*{Tr}(\bar{{\bm{V}}}_{{\bm{A}}_{T}}^{1/2})=O(GT^{1/2}),$

### v2

@@ 3.3 Theoretical Analysis
Following previous work (Gupta et al., 2018; Feinberg et al., 2023), we provide a convergence analysis of the proposed algorithm within the online optimization framework (Hazan et al., 2016; Shalev-Shwartz et al., 2012). In online convex optimization setting, a parameter ${\bm{\theta}}_{t}\in\mathcal{K}$ is chosen iteratively, where $\mathcal{K}$ is a convex decision set. After each decision- ${\bm{\theta}}_{t}$, a convex loss function $f_{t}$ is revealed, potentially chosen adversarially. The regret accumulated by the algorithm up to step $T$ is defined as
@@
Let $\bar{{\bm{X}}}_{{\bm{A}}_{t}}=(\bar{{\bm{V}}}_{{\bm{A}}_{t}}+\rho_{{\bm{A}}_{t}}{\bm{I}})^{-1/2}$ be the unmagnified preconditioner ${\bm{P}}_{{\bm{A}}_{t}}=({\bm{U}}_{{\bm{B}}_{t}})^{\top}{\bm{U}}_{{\bm{B}}_{t-1}}$, and ${\bm{Q}}_{{\bm{A}}_{t}}=({\bm{R}}_{{\bm{B}}_{t}}^{\top})^{\dagger}{\bm{R}}_{{\bm{B}}_{t-1}}^{\top}$, then we have
${\|\bar{{\bm{X}}}_{{\bm{A}}_{t}}^{-1}-{\bm{Q}}_{{\bm{A}}_{t}}\bar{{\bm{X}}}_{{\bm{A}}_{t-1}}^{-1}{\bm{Q}}_{{\bm{A}}_{t}}^{T}\|\leq\mu\|\bar{{\bm{X}}}_{{\bm{A}}_{t}}^{-1}-{\bm{P}}_{{\bm{A}}_{t}}\bar{{\bm{X}}}_{{\bm{A}}_{t-1}}^{-1}{\bm{P}}_{{\bm{A}}_{t}}^{T}\|}$
and
$\operatorname*{Tr}({\bm{P}}_{{\bm{A}}_{t}}\bar{{\bm{X}}}_{{\bm{A}}_{t-1}}^{-1}{\bm{P}}_{{\bm{A}}_{t}}^{T})\geq\operatorname*{Tr}(\bar{{\bm{X}}}_{{\bm{A}}_{t-1}}^{-1}+c(\rho_{{\bm{A}}_{t-1}}^{1/2}-\rho_{{\bm{A}}_{t}}^{1/2}){\bm{I}}).$
This assumption essentially constrains the change in ${\bm{U}}_{{\bm{B}}_{t}}$ and ${\bm{R}}_{{\bm{B}}_{t}}$ to be relatively smooth. Under this assumption, we can establish the following stronger convergence result:
@@
${\frac{1}{T}\sum\nolimits_{t=1}^{T}\nabla{\bm{\theta}}_{t}^{\top}({\bm{\theta}}_{t}-{\bm{\theta}}^{*})=O(GD_{\bm{A}}D_{\bm{B}}T^{-1/2}).}$
@@ 4 Related Work
LARS (You et al., 2017) and Lamb (You et al., 2020) are layer-wise adaptive optimization methods originally designed for large batch training. They dynamically adjust the update norm for each weight matrix based on its current norm, which ensures scalar scale invariance. Nonetheless, they still lack transformation invariance.
@@
Gemma-2B Adam 83.76 45.31 24.26 64.0 54.33
LoRA+ 83.75 45.31 23.65 64.4 54.28
ScaledAdam 83.52 45.22 23.96 64.8 54.38
Shampoo 83.26 44.88 23.35 63.6 53.77
Lamb 86.60 47.35 26.76 68.0 57.18
LoRA-RITE 87.28 49.06 30.10 68.8 58.81
@@ 5 Experimental Results
We evaluate the proposed LoRA optimizer against other optimizers across a range of datasets. This includes the Super-Natural Instructions dataset, a comprehensive collection of diverse NLP tasks, as well as four standard LLM benchmarking datasets.
@@
  • Lamb (You et al., 2020): A variant of Adam that normalizes the updates for each layer based on the norm of the parameters.
@@
For each optimizer applied on each data set, we search for the best learning rate from $2*10^{-6}$ to $2*10^{-2}$. The other hyperparameters are listed in the Appendix. For most of the experiments we chose rank $r=16$ for LoRA, based on the ablation study over the rank. We conduct experiments on Gemma (Gemma Team et al., 2024)  2B, 7B, and mT5-XXL (Xue et al., 2021) using TPUs.
@@
We conduct an ablation study on the choice of different LoRA ranks and model architectures. Specifically, we considered rank 4 and 16 on both Gemma 2B (decoder only) and mT5-XXL (encoder-decoder) on the OpenBookQA dataset. As we can see from Table 3, our proposed method performs consistently well across different LoRA ranks. Furthermore, our method can be successfully applied to mT5-XXL which has an encoder-decoder architecture, showing the generalizability of the proposed optimizer.
@@
Table 4: Number of training steps per second for different optimizers. Shampoo preconditioner is updated once every 100 steps. LoRA-RITE has small overhead compared with first-order methods.
@@ 6 Conclusions
Although this work introduces a better optimizer for LoRA, it is important to acknowledge that LoRA itself has limitations. For instance, LoRA has smaller representational power and may result in a minor performance decrease compared to full fine-tuning. Also, how to select rank to strike a good trade-off between efficiency and accuracy may be non-trivial in practice. The work focuses on addressing transformation-invariance when the optimization problem can be written in the form of $f({\bm{A}}{\bm{B}}^{\top})$, and this assumption may not hold for other parameter-efficient structures beyond LoRA. Applying LoRA-RITE to ensure transformation invariance for the other more complicated LoRA variants will be an interesting future direction.
Acknowledgments
We thank Sashank Reddi and Corinna Cortes for useful feedback. The code for our project is available at https://github.com/gkevinyen5418/LoRA-RITE.
@@ A.3 Proof of Theorem 1
Let $\|{\bm{A}}_{1}\|={\bm{\theta}}(n^{a})$, $\|{\bm{B}}_{1}\|={\bm{\theta}}(n^{b})$, $\|\nabla{\bm{Z}}\|={\bm{\theta}}(n^{c})$, $\eta={\bm{\theta}}(n^{d})$, where $\eta$ is the learning rate and $n$ is the network width. Since ${\bm{Z}}={\bm{A}}_{1}{\bm{B}}_{1}^{T}$, from chain rule we know $\nabla{\bm{A}}=\nabla{\bm{Z}}{\bm{B}}$ and $\nabla{\bm{B}}=\nabla{\bm{Z}}^{T}{\bm{A}}$. Since the update rule is symmetric, we can express the updates as
@@
where ${\bm{x}}$ is the input vector, by selecting a proper learning rate $\eta$.
@@ A.4 Proof of Theorem 2
In Algorithm 1, for two equivalent LoRA pairs $({\bm{A}}_{1},{\bm{B}}_{1}),({\bm{A}}_{2},{\bm{B}}_{2})$, their obtained basis ${\bm{U}}_{{\bm{B}}_{1}}$ and ${\bm{U}}_{{\bm{B}}_{2}}$ could be different.
Consequently, we introduce the concept of consistency, which means a matrix output of $f({\bm{A}},{\bm{B}})$ is the same across all equivalent LoRA pairs, up to the choice of the basis.
Definition 3 (Consistency).
For matrix ${\bm{X}}_{{\bm{A}}}\in\mathbb{R}^{m\times r}$, we call it consistent if
${\bm{X}}_{{\bm{A}}_{1}}{\bm{U}}_{{\bm{B}}_{1}}^{\top}={\bm{X}}_{{\bm{A}}_{2}}{\bm{U}}_{{\bm{B}}_{2}}^{\top}\in\mathbb{R}^{m\times n}$
for all equivalent LoRA pairs $({\bm{A}}_{1},{\bm{B}}_{1}),({\bm{A}}_{2},{\bm{B}}_{2})$. Similarly, for ${\bm{H}}_{{\bm{A}}}\in\mathbb{R}^{r\times r}$, we call it consistent if
${\bm{U}}_{{\bm{B}}_{1}}{\bm{H}}_{{\bm{A}}_{1}}{\bm{U}}_{{\bm{B}}_{1}}^{\top}={\bm{U}}_{{\bm{B}}_{2}}{\bm{H}}_{{\bm{A}}_{2}}{\bm{U}}_{{\bm{B}}_{2}}^{\top}\in\mathbb{R}^{n\times n}$
for all equivalent LoRA pairs $({\bm{A}}_{1},{\bm{B}}_{1}),({\bm{A}}_{2},{\bm{B}}_{2})$.
@@
Additionally, to simplify the notation, from now on we use ${\bm{A}},{\bm{B}}$ to stand for ${\bm{A}}_{1},{\bm{B}}_{1}$ and use $\hat{{\bm{A}}},\hat{{\bm{B}}}$ to stand for ${\bm{A}}_{2},{\bm{B}}_{2}$.
Then, we proceed to prove Theorem 2. First, one should note the fact that
${\bm{U}}_{{\bm{B}}}{\bm{U}}_{{\bm{B}}}^{\top}={\bm{U}}_{\hat{{\bm{B}}}}{\bm{U}}_{\hat{{\bm{B}}}}^{\top}$
for all equivalent LoRA pairs. Consequently,
${\bm{U}}_{{\bm{B}}}(\bar{\nabla}{{\bm{A}}})^{\top}\bar{\nabla}{{\bm{A}}}{\bm{U}}_{{\bm{B}}}^{\top}={\bm{U}}_{{\bm{B}}}{\bm{U}}_{{\bm{B}}}^{\top}\nabla{\bm{Z}}^{\top}\nabla{\bm{Z}}{\bm{U}}_{{\bm{B}}}{\bm{U}}_{{\bm{B}}}^{\top}$ (21)
$=$ ${\bm{U}}_{\hat{{\bm{B}}}}{\bm{U}}_{\hat{{\bm{B}}}}^{\top}\nabla{\bm{Z}}^{\top}\nabla{\bm{Z}}{\bm{U}}_{\hat{{\bm{B}}}}{\bm{U}}_{\hat{{\bm{B}}}}^{\top}={\bm{U}}_{\hat{{\bm{B}}}}(\bar{\nabla}{{\bm{A}}})^{\top}\bar{\nabla}{{\bm{A}}}{\bm{U}}_{\hat{{\bm{B}}}}^{\top},$
which shows that (¯$(\bar{\nabla}{{\bm{A}}})^{\top}\bar{\nabla}{{\bm{A}}}$ is consistent.
Second, we can prove $\bar{{\bm{V}}}_{{\bm{A}}_{t}}$ is consistent by mathematical induction. The base case $\bar{{\bm{V}}}_{{\bm{A}}_{0}}=\bf{0}$ is consistent. Assuming $\bar{{\bm{V}}}_{{\bm{A}}_{t-1}}$ is consistent, then
[…]
which means ${{\bm{P}}_{{\bm{A}}_{t}}}\bar{{\bm{V}}}_{{\bm{A}}_{t-1}}{\bm{P}}_{{\bm{A}}_{t}}^{\top}$ and thus
$\bar{{\bm{V}}}_{{\bm{A}}_{t}}={{\bm{P}}_{{\bm{A}}_{t}}}\bar{{\bm{V}}}_{{\bm{A}}_{t-1}}{\bm{P}}_{{\bm{A}}_{t}}^{\top}+\bar{\nabla}{{\bm{A}}_{t}}^{\top}\bar{\nabla}{{\bm{A}}_{t}}$
is consistent.
@@
we have
$\bar{{\bm{S}}}_{{\bm{A}}_{t}}{\bm{U}}_{{\bm{B}}_{t}}^{\top}=\bar{\nabla}{{\bm{A}}_{t}}(\bar{{\bm{V}}}_{{\bm{A}}_{t}}+\rho_{{\bm{A}}_{t}}{\bm{I}})^{-1/2}{\bm{U}}_{{\bm{B}}_{t}}^{\top}$
$=$ ¯$\bar{\nabla}{{{\bm{Z}}_{t}}}{\bm{U}}_{{\bm{B}}_{t}}(\bar{{\bm{V}}}_{{\bm{A}}_{t}}+\rho_{{\bm{A}}_{t}}{\bm{I}})^{-1/2}{\bm{U}}_{{\bm{B}}_{t}}^{\top}$
$=$ ¯$\bar{\nabla}{{{\bm{Z}}_{t}}}({\bm{U}}_{{\bm{B}}_{t}}\bar{{\bm{V}}}_{{\bm{A}}_{t}}{\bm{U}}_{{\bm{B}}_{t}}^{\top}+\rho_{{\bm{A}}_{t}}{\bm{U}}_{{\bm{B}}_{t}}{\bm{U}}_{{\bm{B}}_{t}}^{\top})^{-1/2}$
$=$ ¯$\bar{\nabla}{{{\bm{Z}}_{t}}}({\bm{U}}_{\hat{{\bm{B}}_{t}}}\bar{{\bm{V}}}_{\hat{{\bm{A}}_{t}}}{\bm{U}}_{\hat{{\bm{B}}_{t}}}^{\top}+\rho_{\hat{{\bm{A}}_{t}}}{\bm{U}}_{\hat{{\bm{B}}_{t}}}{\bm{U}}_{\hat{{\bm{B}}_{t}}}^{\top})^{-1/2}$
$=$ $\bar{{\bm{S}}}_{\hat{{\bm{A}}_{t}}}{\bm{U}}_{\hat{{\bm{B}}_{t}}}^{\top}.$
Thus, we can similarly proof by induction that both $\bar{{\bm{S}}}_{{\bm{A}}_{t}}$ and $\bar{{\bm{M}}}_{{\bm{A}}_{t}}$ are consistent, which completes our proof.
@@ A.5 Proof of Theorem 3
$\bar{{\bm{X}}}_{{\bm{A}}_{t}}\equiv(\bar{{\bm{V}}}_{{\bm{A}}_{t}}+\rho_{{\bm{A}}_{t}}{\bm{I}})^{-1/2}$
Then for the ${\bm{A}}$ factor, we have
$\sum_{t=1}^{T}\operatorname{vec}(\nabla{\bm{A}}_{t})^{\top}\operatorname{vec}(\delta{\bm{A}}_{t})=\sum_{t=1}^{T}\operatorname*{Tr}(\nabla{\bm{A}}_{t}^{\top}\delta{\bm{A}}_{t})=\eta\sum_{t=1}^{T}\|\nabla{\bm{A}}_{t}\|_{{\bm{X}}_{{\bm{A}}_{t}}}^{2}$ (23)
$=$ $\eta\sum_{t=1}^{T}\|\bar{\nabla}{\bm{A}}_{t}\|_{\bar{{\bm{X}}}_{{\bm{A}}_{t}}}^{2}\leq 2\eta\operatorname*{Tr}((\sum_{i=1}^{T}\bar{\nabla}{\bm{A}}_{i}^{\top}\bar{\nabla}{\bm{A}}_{i})^{1/2})=O(GT^{1/2}),$
where the inequality comes from Lemma 2 and the fact that
$\bar{{\bm{X}}}_{{\bm{A}}_{t}}=(\bar{{\bm{V}}}_{{\bm{A}}_{t}}+\rho_{{\bm{A}}_{t}}{\bm{I}})^{-1/2}\preceq(\sum_{i=1}^{t}\bar{\nabla}{\bm{A}}_{i}^{\top}\bar{\nabla}{\bm{A}}_{i})^{-1/2}.$
@@
A.6 Proof of Theorem 4
To prove Theorem 4, we start by bounding the escape mass.
Lemma 3.
For the escaped mass $\rho_{{\bm{A}}_{T}}$ in Algorithm 1, we have
$\rho_{{\bm{A}}_{T}}\leq\sum_{t=1}^{T}\operatorname*{Tr}((\bar{\nabla}{{\bm{A}}_{t}})^{\top}\bar{\nabla}{{\bm{A}}_{t}})-\operatorname*{Tr}(\bar{{\bm{V}}}_{{\bm{A}}_{T}})=O(G^{2}T).$
Proof of Lemma 3
Since ${\bm{P}}_{{\bm{A}}_{t}}{\bm{P}}_{{\bm{A}}_{t}}^{\top}\preceq{\bm{I}}$, by Ostrowski’s inequality we have
$\lambda_{i}(\bar{{\bm{V}}}_{{\bm{A}}_{t-1}})\geq\lambda_{i}({{\bm{P}}_{{\bm{A}}_{t}}}\bar{{\bm{V}}}_{{\bm{A}}_{t-1}}{\bm{P}}_{{\bm{A}}_{t}}^{\top})$
for all $i$. Consequently, we have
$d_{\lambda}(\bar{{\bm{V}}}_{{\bm{A}}_{t-1}},{{\bm{P}}_{{\bm{A}}_{t}}}\bar{{\bm{V}}}_{{\bm{A}}_{t-1}}{\bm{P}}_{{\bm{A}}_{t}}^{\top})\leq\operatorname*{Tr}(\bar{{\bm{V}}}_{{\bm{A}}_{t-1}})-\operatorname*{Tr}({{\bm{P}}_{{\bm{A}}_{t}}}\bar{{\bm{V}}}_{{\bm{A}}_{t-1}}{\bm{P}}_{{\bm{A}}_{t}}^{\top}).$
Thus,
$\rho_{{\bm{A}}_{T}}\leq\sum_{t=1}^{T}\operatorname*{Tr}(\bar{{\bm{V}}}_{{\bm{A}}_{t-1}})-\sum_{t=1}^{T}\operatorname*{Tr}({{\bm{P}}_{{\bm{A}}_{t}}}\bar{{\bm{V}}}_{{\bm{A}}_{t-1}}{\bm{P}}_{{\bm{A}}_{t}}^{\top})$
$=$ $\sum_{t=1}^{T}\operatorname*{Tr}((\bar{\nabla}{{\bm{A}}_{t}})^{\top}\bar{\nabla}{{\bm{A}}_{t}})-\operatorname*{Tr}(\bar{{\bm{V}}}_{{\bm{A}}_{T}})=O(G^{2}T).$